# Paprapee Buason, Sidhant Misra Los Alamos National Laboratory Los Alamos, NM, USA

### v2

@@ III-B Error function
Figure 3: An example of an error function where $g(\epsilon)=\epsilon$ and $h(\epsilon)=e^{2\epsilon}-1$.
In contrast, our CBLA approach offers the flexibility to configure an error function that accommodates violations for specific samples, all while considering a predefined cost associated with these violations. This ability to tailor the error function empowers us to strike a balance between accuracy and the acceptable level of conservativeness for the specific system and operating range of interest. In Fig. 3, an example of an error function $f(\epsilon)$ for an overestimating CBLA is depicted. The function $f(\epsilon)$ exhibits a high value for $\epsilon>0$, $f(\epsilon)$ maintains a relatively low value for $\epsilon<0$, and $f(0)$ is zero, indicating an exact approximation.
In this setup, assigning a higher penalty for violations of conservativeness leads to more conservative approximations at the expense of some accuracy. Additionally, when the derivative of the function $h(\epsilon)$ increases, as shown in Fig. 3, the function tends to allow only small positive values of $\epsilon$, because larger positive values result in exponentially higher penalties. Conversely, reducing the penalty cost tends to produce more accurate approximations but with less conservativeness.
@@
When the error function is defined as a piecewise linear function (i.e., when both $g(\epsilon)$ and $h(\epsilon)$ are linear functions), the problem formulation to compute CBLAs in (5)–(7) can be framed as a linear program as follows:
@@
[…]
where $z$ is a slack variable, and $k_{1}$ and $k_{2}$ are the coefficients of the linear error functions, i.e., $g(\epsilon)=k_{1}\epsilon$ and $h(\epsilon)=k_{2}\epsilon$.
Nonlinear error functions transform the regression problem into a mixed-integer nonlinear program. This can be efficiently implemented using a user-defined function in Julia and solved with packages such as Optim, which offers a robust framework for constrained optimization problems [20].
@@ IV Unit Commitment
The variability and uncertainty associated with loads, particularly when load shedding may be necessary, and generation introduce additional complexities that the unit commitment (UC) problem must address to ensure grid reliability and operational efficiency. As a mixed-integer optimization problem, the UC formulation involves binary variables to represent the on/off status of generators, along with continuous variables to model power output.
@@ IV-A Computational procedure and parallelization potential
Figure 4: A flowchart illustrating the computational process for solving the unit commitment (UC) problem using CBLA. Steps enclosed in red boxes correspond to CBLA, while steps enclosed in blue boxes relate to the UC problem. Steps marked with * indicate parallelizable processes.
Solving the unit commitment problem using CBLA involves multiple computational steps, each contributing to accurately determining generator commitments and dispatch while leveraging efficient approximations of power flow equations. Given the complexity of unit commitment and the nonlinearity of AC power flow, CBLA serves as an effective method to construct computationally efficient linear approximations. The overall process consists of three key steps:
@@
    Unit commitment optimization: Using the CBLA approximations obtained in Step 1, solve the UC problem to determine generator statuses for each generator at each hour while ensuring operational constraints are satisfied.
@@
    AC optimal power flow (AC OPF) evaluation: Given the generator statuses from Step 2, solve the AC OPF that allows load shedding to determine the generation dispatch at each hour, ensuring feasibility under AC power flow constraints [21].
This computational framework allows parallel execution, particularly in Steps 1 and 3. In Step 1, solving CBLA involves two key computations that can be parallelized. First, the power flow solutions at each operating point can be computed independently. Second, the regression problem for the CBLA can be solved independently for each quantity, using the same set of samples across all regression problems. Similarly, in Step 3, since ramping constraints are not enforced and only minimum up- and down-time constraints are considered, the AC OPF computations at different time steps are independent once generator statuses are fixed, enabling parallel execution across time periods. Fig. 4 illustrates the full set of computational steps required to solve the unit commitment problem using the CBLA approach.
@@ IV-B Optimization Problem Formulation
Let $u_{g,t}$ be a generator on/off status and $s_{g,t}\;(r_{g,t})$ be a generator startup (shutdown) status for generator $g$ at time $t$. The unit commitment problem with the DC power flow [22] is formulated as follows:
$\min$ $\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+C_{g}^{\text{startup}}s_{g,t}+C_{g}^{\text{shutdown}}r_{g,t}\right)$ (10a)
@@ IV-C Unit commitment using conservative bias linear approximations
The unit commitment problem formulated in the previous subsection enforces network feasibility using a DC power flow approximation. While computationally efficient, the DC power flow model does not explicitly incorporate voltage and current constraints. In this section, we introduce penalty terms and replace the DC power flow constraints in (10g)–(10h) with a CBLA, which provides a linearized yet conservative representation of these constraints.
@@
To maintain the validity of CBLA-based constraints during the UC process, we adopt an iterative approach (Fig. 4). In each iteration, we solve the UC problem with the current CBLA, then check if the generation statuses and dispatch levels match those from a first-order Taylor-based method or previous CBLA iterations. We also verify whether power injections stay within the sampled region used to build the CBLA. If violations occur, we identify the affected time periods and buses or lines, draw new samples around the updated solution, recompute the CBLA, and resolve the UC problem. This process prevents accuracy degradation from extrapolation and cycling between inconsistent UC solutions.
This CBLA-based formulation further allows two practical enhancements. First, it offers flexibility in controlling conservativeness by adjusting a bias term when constructing the constraints, enabling the user to trade off between feasibility margin and conservatism. Second, it supports the inclusion of penalty terms that allow for controlled constraint violations, enhancing feasibility in cases where a strictly conservative formulation may be overly restrictive.
@@ IV-D Combinatorial Growth of Valid Generator Statuses
In the unit commitment formulation, each generator’s on/off status over time must satisfy minimum up and down time constraints (10e)–(10f), substantially reducing the number of admissible on/off sequences. Nevertheless, the feasible status space still grows exponentially with the number of time periods, making brute-force enumeration impractical.
@@
$f(\lvert\mathcal{T}\rvert,s,\ell)=\begin{cases}\sum\limits_{k=d}^{\lfloor\nicefrac{{\lvert\mathcal{T}\rvert}}{{2}}\rfloor}f(\lvert\mathcal{T}\rvert-k,1-s,k),\;\text{if }\ell=d,\lvert\mathcal{T}\rvert\geq d\\[4.30554pt] f(\lvert\mathcal{T}\rvert-1,s,\ell-1),\;\text{if }\ell>d\\[4.30554pt] 1,\;\text{if }\lvert\mathcal{T}\rvert=\ell\leq d\text{, one of the exceptions holds}\\[4.30554pt] 0,\;\text{otherwise}\end{cases}$
@@
For example, with $\lvert\mathcal{T}\rvert=5$ and $d=2$, applying the recurrence yields 13 valid sequences for a single generator. Therefore, for 6 generators, the total number of configurations is $(F(5))^{6}=13^{6}=4,\!826,\!809$, demonstrating how the feasible status space grows rapidly even for modest systems. This growth illustrates why brute-force enumeration of all valid generator commitment combinations quickly becomes intractable. To manage this complexity, we avoid exhaustive search by solving the unit commitment problem iteratively: each iteration only considers a manageable subset of decisions, and additional constraints are introduced in subsequent iterations to guide the solution away from unstable commitment patterns, such as cycling or oscillating generator decisions, i.e., repeatedly looping through statuses. The algorithm is explained in more detail in Section IV-A.
@@ IV-E AC Optimal Power Flow with Load Shedding
The AC Optimal Power Flow (AC OPF) problem determines the optimal generation dispatch by solving the full nonlinear AC power flow equations while satisfying network constraints. Using generator commitment statuses from the unit commitment problem in (10) (solved with formulations such as DC power flow, first-order Taylor approximations, or CBLAs), we compute the dispatch for each online generator at each hour. The multi-period formulation minimizes the total cost across all hours while ensuring power balance and network feasibility at each time. If constraints prevent meeting full demand, load shedding is used as a last resort, with $L_{i,t}$ denoting the load shed at bus $i$ and time $t$. The AC OPF with load shedding is formulated as follows:
$\min$ $\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+\sum_{i\in\mathcal{B}}C^{\text{shed}}(L_{i,t})\right)$ (14a)
@@
where $\beta_{i}=\tan\alpha_{i}$, given that $\cos\alpha_{i}$ represents the power factor associated with load $i$. The objective function in (14a) minimizes the total cost of generation over all time periods while penalizing load shedding using a function $C^{\text{shed}}(L_{i,t})$ to discourage demand curtailment unless necessary. The power balance constraints in (14b)–(14c) enforce the nonlinear AC power flow equations for active and reactive power at each bus and time period. The generator limits in (14d)–(14e) ensure that active and reactive power generation respects the unit commitment status $u_{g,t}$. The voltage limits in (14f) ensure that bus voltages remain within safe operating ranges. The thermal constraints in (14g) restrict apparent power flows on transmission lines to their rated capacities. Finally, the load shedding constraints in (14h) ensure that the curtailed demand remains non-negative and does not exceed the original load at any bus and time period. This formulation integrates the unit commitment results into AC OPF, optimizing the dispatch of committed generators while allowing for load shedding with constant power factors.
@@ V Numerical results
The test cases used in the simulations are the IEEE 24-bus system and the IEEE 30-bus system, all of which are accessible in Matpower [23]. For approximations of voltage and current flow, we draw $500$ samples by varying the power injections within a range of $70\%$ to $130\%$ of their nominal values. Both voltage and current flow values are reported in per unit (pu). We use the $\ell_{1}$ norm as the loss function $\mathcal{L}(\,\cdot\,)$. The numerical simulation was conducted in MATLAB using Yalmip [24] and in Julia using the Optim package.
@@ V-B Application: Unit commitment
The comparison is conducted on the IEEE 30-bus test system, which has 6 generators. Voltage magnitude limits are set to 0.94 p.u. and 1.05 p.u. at all buses, and current magnitude limits are derived from the values in mpc.branch(:,RATE_A) in the MATPOWER case files. The cost of load shedding is $C^{\text{shed}}(L_{i,t})=500L_{i,t}$. We draw $1000$ samples by varying power injections, both loads and generations, to cover possible load shedding and generation variation within a range of $70\%$ to $130\%$ of their nominal values. As shown in Table II, we report both the total cost, which includes load shedding penalties, and the generation-only cost, which excludes them to better isolate the impact of load shedding. Although the DC-based formulation often yields lower generation-only costs, it leads to much higher total costs due to frequent and severe load shedding. As illustrated in Fig. 7, the generator commitment schedules from CBLA differ from those of the DC and first-order Taylor methods, highlighting CBLA’s ability to adapt commitment decisions based on a more accurate power flow approximation.
@@
This trade-off underscores the need to balance conservativeness and approximation accuracy. CBLA’s tunable parameter $\alpha$ allows us to adjust this balance to achieve the desired feasibility and flexibility. By tailoring the approximation to system conditions, CBLA helps reduce constraint violations and yields UC solutions that generally remain feasible under the nonlinear AC power flow model.
Moreover, the benefits of CBLA become more pronounced in larger systems or under tight feasibility margins. Unlike the first-order Taylor approximation, which is inherently local, CBLA is constructed to remain conservative across a sampled region of interest. This range-aware property improves cost-efficiency in decision-making. Compared to traditional methods, CBLA provides a more adaptive solution framework that maintains the tractability of linear programming while reducing the gap between the models of linear and nonlinear power systems.
@@ VI Conclusion and future work
This paper presents a conservative bias linear approximation (CBLA) approach for approximating the power flow equations, aiming to balance conservativeness and accuracy while preserving linearity. The numerical results highlight the potential advantages of using CBLA for power flow problems. Selecting suitable error functions enables an effective balance between conservativeness and accuracy. Additionally, the ability to choose different error functions allows CBLA to be tailored to specific systems and operational conditions. When applied to the unit commitment problem, the CBLA-based formulation consistently outperformed formulations based on DC power flow and first-order Taylor approximations, achieving lower overall operating costs in the subsequent evaluation of the AC optimal power flow.
In our future work, we aim to extend our current approach by developing additional conservative bias approximations through the use of piecewise linearizations. Moreover, we plan to apply our proposed approach to a broader range of power system planning and resilience tasks. This includes tackling complex bilevel problems, conducting capacity expansion planning studies, and extending the application of CBLA to unbalanced three-phase systems.

### v3

@@ III-B Error Function Design
Figure 3: Example CBLA error function for an overestimating approximation, where $g(\epsilon)=\epsilon$ and $h(\epsilon)=e^{2\epsilon}-1$.
An example is shown in Fig. 3, where $g(\epsilon)=\epsilon$ and $h(\epsilon)=e^{2\epsilon}-1$. In this case, positive errors incur exponentially increasing penalties, discouraging large violations while allowing small deviations when necessary. Increasing the growth rate of $h(\epsilon)$ promotes stronger conservativeness, whereas reducing its slope yields a more accurate but less conservative approximation.
@@
When both $g(\epsilon)$ and $h(\epsilon)$ are linear functions, i.e., the error function is piecewise linear, the CBLA regression problem can be reformulated as a linear program:
@@
$z_{m}\geq k_{1}\epsilon_{m},$ (9c)
$z_{m}\geq k_{2}\epsilon_{m},$ (9d)
where $z_{m}$ is an auxiliary variable and $k_{1}$, $k_{2}$ define the slopes of $g(\epsilon)=k_{1}\epsilon$ and $h(\epsilon)=k_{2}\epsilon$, respectively.
If nonlinear penalty functions are selected, the resulting optimization problem becomes a nonlinear program (or potentially a mixed-integer nonlinear program, depending on the formulation). Such problems can be implemented using user-defined functions in Julia and solved using packages such as Optim, which provides a flexible framework for constrained optimization [optimjl].
@@ IV Unit Commitment
The variability and uncertainty associated with loads, particularly when load shedding may be necessary, and generation introduce additional complexities that the UC problem must address to ensure grid reliability and operational efficiency. As a mixed-integer optimization problem, the UC formulation involves binary variables to represent the on/off status of generators, along with continuous variables to model power output.
@@ IV-A Computational Procedure and Parallelization Potential
Figure 4: A flowchart illustrating the computational process for solving the UC problem using CBLA. Steps enclosed in red boxes correspond to CBLA, while steps enclosed in blue boxes relate to the UC problem. Steps marked with * indicate parallelizable processes.
Solving the UC problem using CBLA involves multiple computational steps, each contributing to accurately determining generator commitments and dispatch while leveraging efficient approximations of power flow equations. Given the complexity of UC and the nonlinearity of AC power flow, CBLA serves as an effective method to construct computationally efficient linear approximations. The overall process consists of three key steps:
@@
    UC optimization: Using the CBLA approximations obtained in Step 1, solve the UC problem to determine generator statuses for each generator at each hour while ensuring operational constraints are satisfied.
@@
    AC optimal power flow (AC OPF) evaluation: Given the generator statuses from Step 2, solve the AC OPF that allows load shedding to determine the generation dispatch at each hour, ensuring feasibility under AC power flow constraints [Happ1977].
This computational framework allows parallel execution, particularly in Steps 1 and 3. In Step 1, solving CBLA involves two key computations that can be parallelized. First, the power flow solutions at each operating point can be computed independently. Second, the regression problem for the CBLA can be solved independently for each quantity, using the same set of samples across all regression problems. Similarly, in Step 3, since ramping constraints are not enforced and only minimum up- and down-time constraints are considered, the AC OPF computations at different time steps are independent once generator statuses are fixed, enabling parallel execution across time periods. Fig. 4 illustrates the full set of computational steps required to solve the UC problem using the CBLA approach.
@@ IV-B Optimization Problem Formulation
Let $u_{g,t}$ be a generator on/off status and $s_{g,t}\;(r_{g,t})$ be a generator startup (shutdown) status for generator $g$ at time $t$. The UC problem with the DC power flow [vandenbergh2014] is formulated as follows:
$\min\$$\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+C_{g}^{\text{startup}}s_{g,t}+C_{g}^{\text{shutdown}}r_{g,t}\right)$ (10a)
@@ IV-C Unit Commitment using Conservative Bias Linear Approximations
The UC problem formulated in the previous subsection enforces network feasibility using a DC power flow approximation. While computationally efficient, the DC power flow model does not explicitly incorporate voltage and current constraints. In this section, we introduce penalty terms and replace the DC power flow constraints in (10g)–(10h) with a CBLA, which provides a linearized yet conservative representation of these constraints.
@@
To maintain the validity of CBLA-based constraints during the UC process, we adopt an iterative approach (Fig. 4). In each iteration, the UC problem is solved using the current CBLA, and the resulting generation statuses and dispatch levels are checked against those from a first-order Taylor-based method or previous CBLA iterations. We also verify that power injections remain within the sampled region used to construct the CBLA. If violations are detected, new samples are drawn around the updated solution, the CBLA is recomputed, and the UC problem is resolved. This process mitigates extrapolation errors and prevents cycling between inconsistent UC solutions.
In addition, the search can be restricted to a limited set of candidate UC solutions generated by the solver. For instance, a small number of feasible candidates (e.g., up to 20) may be evaluated and compared based on feasibility and operational cost, with each candidate validated through an AC OPF evaluation. This approach provides a practical trade-off between solution quality and computational effort by selecting the best available candidate within the allowed computation time.
This CBLA-based formulation further allows two practical enhancements. First, it offers flexibility in controlling conservativeness by adjusting a bias term when constructing the constraints, enabling the user to trade off between feasibility margin and conservativeness. Second, it supports the inclusion of penalty terms that allow for controlled constraint violations, enhancing feasibility in cases where a strictly conservative formulation may be overly restrictive.
@@ IV-D Combinatorial Growth of Valid Generator Statuses
In the UC formulation, each generator’s on/off status over time must satisfy minimum up and down time constraints (10e)–(10f), substantially reducing the number of admissible on/off sequences. Nevertheless, the feasible status space still grows exponentially with the number of time periods, making brute-force enumeration impractical.
@@
[…]
@@
For example, with $\lvert\mathcal{T}\rvert=5$ and $d=2$, applying the recurrence yields 13 valid sequences for a single generator. Therefore, for six generators, the total number of configurations is $(F(5))^{6}=13^{6}=4,\!826,\!809$, demonstrating how the feasible status space grows rapidly even for modest systems. This growth illustrates why brute-force enumeration of all valid generator commitment combinations quickly becomes intractable. To manage this complexity, we avoid exhaustive search by solving the UC problem iteratively: each iteration only considers a manageable subset of decisions, and additional constraints are introduced in subsequent iterations to guide the solution away from unstable commitment patterns, such as cycling or oscillating generator decisions, i.e., repeatedly looping through statuses. The algorithm is explained in more detail in Section IV-A.
@@ IV-E AC Optimal Power Flow with Load Shedding
The AC Optimal Power Flow (AC OPF) problem determines the optimal generation dispatch by solving the full nonlinear AC power flow equations while satisfying network constraints. Using generator commitment statuses from the UC problem in (10) (solved with formulations such as DC power flow, first-order Taylor approximations, or CBLAs), we compute the dispatch for each online generator at each hour. The multi-period formulation minimizes the total cost across all hours while ensuring power balance and network feasibility at each time. If constraints prevent meeting full demand, load shedding is used as a last resort, with $L_{i,t}$ denoting the load shed at bus $i$ and time $t$. The AC OPF with load shedding is formulated as follows:
$\min\$$\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+\sum_{i\in\mathcal{B}}C^{\text{shed}}(L_{i,t})\right)$ (14a)
@@
where $\beta_{i}=\tan\alpha_{i}$, given that $\cos\alpha_{i}$ represents the power factor associated with load $i$. The objective function in (14a) minimizes the total cost of generation over all time periods while penalizing load shedding using a function $C^{\text{shed}}(L_{i,t})$ to discourage demand curtailment unless necessary. The power balance constraints in (14b)–(14c) enforce the nonlinear AC power flow equations for active and reactive power at each bus and time period. The generator limits in (14d)–(14e) ensure that active and reactive power generation respects the UC status $u_{g,t}$. The voltage limits in (14f) ensure that bus voltages remain within safe operating ranges. The thermal constraints in (14g) restrict apparent power flows on transmission lines to their rated capacities. Finally, the load shedding constraints in (14h) ensure that the curtailed demand remains non-negative and does not exceed the original load at any bus and time period. This formulation integrates the UC results into AC OPF, optimizing the dispatch of committed generators while allowing for load shedding with constant power factors.
@@ V Numerical results
The test cases used in the simulations are the IEEE 24-bus system and the IEEE 30-bus system, both of which are accessible in Matpower [zimmerman_matpower_2011]. For approximations of voltage and current flow, we draw $500$ samples by varying the power injections within a range of $70\%$ to $130\%$ of their nominal values. Both voltage and current flow values are reported in per unit (pu). We use the $\ell_{1}$ norm as the loss function $\mathcal{L}(\,\cdot\,)$. The numerical simulation was conducted in MATLAB using Yalmip [Lofberg2004] and in Julia using the Optim package.
The number of samples (500–1000 in the reported experiments) is selected to provide adequate coverage of the multidimensional injection space while maintaining computational tractability. Empirically, we observe that beyond this range (e.g., more than 500–1000 samples for the tested systems), the marginal improvement in approximation accuracy and solution robustness becomes negligible relative to the additional computational cost.
@@ V-B Application: Unit Commitment
The comparison is conducted on the IEEE 30-bus test system, which has six generators. Voltage magnitude limits are set to 0.94 p.u. and 1.05 p.u. at all buses, and current magnitude limits are derived from the values in mpc.branch(:,RATE_A) in the Matpower case files. The cost of load shedding is $C^{\text{shed}}(L_{i,t})=500L_{i,t}$. We draw $1000$ samples by varying power injections, both loads and generations, to cover possible load shedding and generation variation within a range of $70\%$ to $130\%$ of their nominal values. As shown in Table II, we report both the total cost, which includes load shedding penalties, and the generation-only cost, which excludes them to better isolate the impact of load shedding. Although the DC-based formulation often yields lower generation-only costs, it leads to much higher total costs due to frequent and severe load shedding. As illustrated in Fig. 7, the generator commitment schedules from CBLA differ from those of the DC and first-order Taylor methods, highlighting CBLA’s ability to adapt commitment decisions based on a more accurate power flow approximation.
@@
This trade-off underscores the need to balance conservativeness and approximation accuracy. CBLA’s tunable parameter $\alpha$ allows for adjusting this balance to achieve the desired level of feasibility robustness and operational flexibility. By tailoring the approximation to system conditions, CBLA helps reduce constraint violations and yields UC solutions that generally remain feasible under the nonlinear AC power flow model.
A practical approach to selecting $\alpha$ is to evaluate constraint violations on an independent set of AC-feasible operating points and adjust the parameter accordingly. If $\alpha$ is chosen too small, the approximation may become insufficiently conservative, potentially leading to AC infeasibility in the UC solution. Conversely, excessively large values of $\alpha$ result in a more conservative formulation, which preserves feasibility but may increase operational cost. In either case, the impact primarily affects the optimality–conservativeness balance rather than the structural validity of the method.
Moreover, the benefits of CBLA become more pronounced in larger systems or under tight feasibility margins. Unlike the first-order Taylor approximation, which is inherently local, CBLA is constructed to remain conservative across a sampled region of interest. This range-aware property improves cost efficiency. Compared to traditional methods, CBLA provides a more adaptive solution framework that maintains the tractability of linear programming while reducing the gap between linear and nonlinear power system models.
@@ VI Conclusion and Future Work
This paper presents a conservative bias linear approximation (CBLA) approach for approximating the power flow equations, aiming to balance conservativeness and accuracy while preserving linearity. The numerical results highlight the potential advantages of using CBLA for power flow problems. Selecting suitable error functions enables an effective balance between conservativeness and accuracy. Additionally, the ability to choose different error functions allows CBLA to be tailored to specific systems and operational conditions. When applied to the UC problem, the CBLA-based formulation consistently outperformed formulations based on DC power flow and first-order Taylor approximations, achieving lower overall operating costs in the subsequent evaluation of the AC optimal power flow. For benchmarking purposes, we compared CBLA against the DC power flow model and the first-order Taylor approximation, as these remain among the most widely adopted linearization approaches in both industry practice and academic research. In particular, DC power flow is the standard approximation for large-scale UC formulations, while Taylor-based linearizations are commonly used in optimization-based power system studies. A systematic comparison with other advanced linearized AC models (e.g., LPAC [Coffrin2014] and other formulations [fnt]) is a topic for future work.
In addition, our future work aims to extend our current approach by developing additional CBLA through the use of piecewise linearizations. Moreover, we plan to apply our proposed approach to a broader range of power system planning and resilience tasks. This includes tackling complex bilevel problems, conducting capacity expansion planning studies, and extending the application of CBLA to unbalanced three-phase systems.